A seed merchant can hire additional labor at $7.50 per hour. Hiring five more hours of labor will increase output by ten units. If the final product sells for $3.50, should extra labor be hired? Explain.
Answer #1

Answer

No,

the firm hires labor if the marginal revenue product of the output is equal or higher than the wage.

the total wage for the next 5 hours =7.5*5=37.5

total revenue from it =10*3.5=$35

the revenue is less than the wage paid so the firm should not hire it.

or

Marginal product per hour =2 unit

MRP=MP*price=3.5*2=$7

the MRP<wage so the firm should not hire it.